# A Comparison of Relativistic Coupled Cluster and Equation of Motion Coupled Cluster Quadratic Response Theory

### v1

@@ 4.2 Verdet constant of Xe, Rn and Og
In table 4 we present our results for the calculation of the Verdet constant for Xe, Rn and Og. In line with our prior work, we considered three laser wavelengths (589 nm, 694.3 nm, and 1064 nm) to compare QR-EOMCC and QR-CC for different Hamiltonians. The Verdet constants are evaluated as quadratic response functions30, 73:
@@
with the number density of the gas, the elementary charge, the electron mass, the speed of light in vacuo, and is magnetic dipole moment operator. Given the atomic symmetry, we only needed to calculate one unique component and could obtain through the expression
@@
Comparing first the differences between QR-CC and QR-EOMCC, we observe that these become larger as the wavelength decreases, and increase as well as the atoms become heavier. However, if we consider the magnitude of these differences, we see that the differences are rather small overall, between 1.1% and 1.3%, and are very similar for the different wavelengths. These additional calculations with respect to those already reported by 115, allow us to more clearly separate contributions from different relativistic effects. With respect to the Hamiltonian and as discussed previously, the Verdet constant for Xe is still well described by non-relativistic Hamiltonians 146, 115, and scalar relativity is the dominant relativistic effect. Furthermore, we observe no difference between (SF-)X2C and (SF-)2DCM results, and find only rather small differences X2C and 2DCM. As this is an atomic system, these differences are primarily due to the difference in Hamiltonian used to compute the spin-orbit screening in the mean-field approximation used in the X2C approach. Adding the Gaunt interaction to this mean-field screening acts to slightly decrease the 2DCM value, in the direction of the spinfree and X2C (Douglas-Kroll-Hess based screening) result. We also observe that in spite of the fact that QR-CC values are smaller than QR-EOMCC ones, they still overestimate the experimental results. For Rn relativistic effects are important and spin-orbit coupling effects significantly increase the Verdet constant. We again see some small differences between X2C and 2DCM. For Og, spin-free results with X2C and 2DCM are roughly the same, indicating that the bare nucleus and screened X2C transformation matrices are very similar (this is the only difference between the two spin-free approaches, see 122 for a more extensive discussion). With spin-orbit coupling differences between X2C and 2DCM are more noticeable at about 3%. For this superheavy element, the spin-orbit effect is very significant and even dominates over scalar relativity, with 2DCM values being roughly three times that of SF-2DCM. While much larger in magnitude, the Gaunt interaction behaves similarly in Og as in Xe, reducing the Verdet constant at all frequencies.
@@
In figure 4 we present the frequency dependence of the Verdet constant obtained with the 2DCM Hamiltonian (see SI for the values for each atom at different frequencies). From these data, we observe that no significant discrepancy arises from the use of QR-EOMCC instead of QR-CC. Across the range of studied wave lenghts, the differences are only of the order of 1.1-1.3 %. The magnitude of the Verdet constant for Rn is found to be roughly twice of that for Xe across the range of studied wavelenghts, while the one of Og is much larger still, with its magnitude being 4.5 times larger for the longest to about 6.5 times larger for the shortest wavelenghts. This is in stark contrast to what is observed for the spin-free calculation presented in table 4, where we see changes of only a factor of roughly 1.5 going from Xe to Rn, and of 1.8 going from Rn to Og. This is consistent with prior theoretical investigations of Og, that predict this element to be much more polarizable due to the extremely large spin-orbit effects in this superheavy element 147, 148. While the first excited state is lowered due to relativity, this strongly increased polarizability due to spin-orbit effects manifests itself already far away from this resonance. To document this, we include in the SI calculations for the lowest excited states of these atoms. These show that while the first excited state is progressively lowered by spin-orbit coupling as the atom gets heavier, the wavelengths probed are still far away from the resonances.
@@
To further complete the picture, we computed the 2DCM dipole polarizabilities for the noble gases with LR-CC and LR-EOMCC for the same range of frequencies as for the Verdet constant. The results, presented in the SI, are in line with the findings for the static polarizabilties for the HX systems and our work 113 in that there is no appreciable difference between the two CC variants for this property. Our results are in good agreement with the experimental static dipole polarizabilities for Xe and Rn (27.3 and 35.8 a.u. respectively) 150, 151, and likewise for the theoretical predictions for Og (57.98 a.u.) of Smits and coworkers 69. Though we did not calculated values for atomic Ts, our results from HTs capture well the trend of decreasing polarizability from Ts to Og (76.3 to 57.98 a.u.) 69 predicted in that work. We attribute the observed discrepancies to the lack of higher excitations in CCSD.
@@
To check for the contribution of subvalence electrons, we also considered additional calculations in which we retained the virtual space truncation at 5 a.u. but included also the electrons. In our prior QR-EOMCC calculations 115 we found, based on non-relativistic calculations (the difference of calculations with the same correlating space, and by correlating all occupied and virtual orbitals), that the effect of correlating electrons beyond the valence shell could result in a decrease of the Verdet constant by about -0.08 [10-3 rad/(T m)] for Xe. Our current results, shown in the SI, confirm that this trend is indeed followed by all atoms, though with the current virtual space truncation we only recover about half of this value. This change is consistent with our earlier finding, that upon correlating the electrons, the excitation energies for all systems are (slightly) shifted upward. Another change seen upon correlating the electrons is that the difference between QR-CC and QR-EOMCC values increases somewhat, though in relative terms these differences remain close to 1% overall.
To conclude this section we note that, unlike the hyperpolarizability case, our results indicate that QR-CC and QR-EOMCC perform very much alike for the Verdet constant, and notably better than common density functional approximations such as B3LYP 115. To keep computational costs manageable compared to such more approximate, but also much more efficient approaches, it may indeed be advantageous to resort to QR-EOMCC given its somewhat smaller memory footprint and cost as compared to QR-CC.
@@ 5 Conclusions
In this work we have presented an implementation in the GPU-accelerated ExaCorr module of coupled cluster quadratic response (QR-CC) theory which can be used in combination with both (2- and 4-component based) relativistic and non-relativistic Hamiltonians, thus extending our prior implementation quadratic response based on the equation of motion (QR-EOMCC) formalism 115.
@@
We have showcased this implementation by carrying out a detailed comparison between QR-CC and QR-EOMCC in the calculation of both static and dynamic first hyperpolarizabilities for the hydrogen halide series (for HF to HTs), as well as the calculation of the Verdet constant for noble gases (Xe, Rn and Og). In addition to the comparison of non-relativistic and X2C Hamiltonians as done previously, we have carried out calculations with the X2C molecular mean field approach, and gauged the contribution of spin-orbit coupled and scalar relativistic effects.
For the first hyperpolarizabilities, we find that for the HF molecule the difference between QR-CC and EOM-CC is relatively small (around 4%), in line with the studies reported in the literature employing non-relativistic approaches that do not see significant differences between the two approaches. The differences between QR-CC and QR-EOMCC do however increase as one goes down the periodic table, with differences of about 20% already for HCl and a sign change for in HI. The overall shape of the dispersion curves remains unchanged, indicating that the EOM approximation yields qualitatively but not always quantitatively correct results. The contribution of spin-orbit coupling to the hyperpolarizability is rather small up to HBr, but becomes significant from HI onwards, also for the static case. Truncation of the correlating space appears to affect QR-EOMCC more than QR-CC for the component for HBr, but the additional electron correlation recovered does not modify the trends upon going down the series from HF to HBr.
@@
In contrast to the hyperpolarizabilies, for the Verdet constant the difference between QR-CC and EOM-CC is minor and roughly constant percentage-wise (about 1%) for all three systems studied. Spin-orbit coupling accounts for a relatively small part of the response (about 4%) in Xe, becomes important for Rn (about 35-40%) and absolutely dominates for Og. Correlating the d electrons in addition to the valence reduces the magnitude of the Verdet constants by about 10%.
@@
The calculations presented here, albeit restricted to a relatively small set of atomic and molecular systems serve to highlight similarities, but also sometimes important differences, in behavior between the QR-CC and QR-EOMCC approaches across the periodic table and for different molecular properties. While QR-EOMCC is advantageous in terms of computational cost and implementation complexity with respect to QR-CC, it is thus important to further characterize differences between the two approaches in the calculation of nonlinear properties for heavier element systems.
@@
This research used resources of the Oak Ridge Leadership Computing Facility, which is a DOE Office of Science User Facility supported under Contract DE-AC05-00OR22725 (allocations CHM160, CHM191 and CHP109). XY, LH and ASPG acknowledge funding from projects CPER WaveTech, Labex CaPPA (Grant No. ANR-11-LABX-0005-01), ANR CompRIXS (Grant Nos. ANR-19-CE29-0019 and DFG JA 2329/6-1), ANR SCREECHES (Grant Nos. ANR-24-CE29-0904) the I-SITE ULNE project OVERSEE and MESONM International Associated Laboratory (LAI) (Grant No. ANR-16-IDEX-0004), as well support from the French national supercomputing facilities (Grant Nos. DARI A0150801859, A0170801859) and the HPC center at the University of Lille. For the purpose of Open Access, a CC-BY public copyright license has been applied by the authors to the present document and will be applied to all subsequent versions up to the Author Accepted Manuscript arising from this submission.
@@
The data (input/output) corresponding to the calculations of this paper are available at the Zenodo repository under DOI: 10.5281/zenodo.15738821.
The Supporting Information is available free of charge on the ACS Publications website at DOI: XXX.

### v3

@@ 4.2 Verdet constant of noble gases (Xe-Og) and hydrogen halides (HF-HI)
In Table 5 we present our results for the calculation of the Verdet constant for Xe, Rn and Og. In line with our prior work, we considered three laser wavelengths (589.0 nm, 694.3 nm, and 1064.0 nm) to compare QR-EOMCC and QR-CC for different Hamiltonians. The Verdet constants are evaluated as quadratic response functionscoriani_coupled_1997, norman_quadratic_2004:
@@
with the number density of the gas, the elementary charge, the electron mass, the speed of light in vacuo, and is magnetic dipole moment operator. For the noble gases, due to atomic symmetry, we only needed to calculate one unique component and could obtain through the expression
@@
Comparing first the differences between QR-CC and QR-EOMCC, we observe that these become larger as the wavelength decreases, and increase as the atoms become heavier. However, if we consider the magnitude of these differences, we see that the differences are rather small overall (about 1 to 1.4%). Furthermore, the differences are very similar for the different wavelengths. These additional calculations with respect to those already reported by Yuan_Xiang_QR, allow us to more clearly separate contributions from different relativistic effects. With respect to the Hamiltonian and as discussed previously, the Verdet constant for Xe is still well described by non-relativistic Hamiltonians cadene_circular_2015, Yuan_Xiang_QR, and scalar relativity is the dominant relativistic effect. Furthermore, we observe no difference between (SF-)X2C and (SF-)2DCM results, and find only rather small differences X2C and 2DCM. As this is an atomic system, these differences are primarily due to the difference in Hamiltonian used to compute the spin-orbit screening in the mean-field approximation used in the X2C approach. Adding the Gaunt interaction to this mean-field screening acts to slightly decrease the 2DCM value, in the direction of the spinfree and X2C (Douglas-Kroll-Hess based screening) result. We also observe that in spite of the fact that QR-CC values are smaller than QR-EOMCC ones, they still overestimate the experimental results. For Rn relativistic effects are important and spin-orbit coupling effects significantly increase the Verdet constant. We again see some small differences between X2C and 2DCM. For Og, spin-free results with X2C and 2DCM are roughly the same, indicating that the bare nucleus and screened X2C transformation matrices are very similar (this is the only difference between the two spin-free approaches, see sikkema2009molecular for a more extensive discussion). With spin-orbit coupling differences between X2C and 2DCM are more noticeable at about 3%. For this superheavy element, the spin-orbit effect is very significant and even dominates over scalar relativity, with 2DCM values being roughly three times that of SF-2DCM. While much larger in magnitude, the Gaunt interaction behaves similarly in Og as in Xe, reducing the Verdet constant at all frequencies.
In Figure 4 and LABEL:SI-tab:verdet-table-scan-nsnp we present the frequency dependence of the Verdet constant obtained with the 2DCM Hamiltonian for additional frequencies than those in Table 5. From these data, we observe that no significant discrepancy arises from the use of QR-EOMCC instead of QR-CC. Across the range of studied wave lengths, the differences remain on the order of 1.1-1.3%.
@@
The magnitude of the Verdet constant for Rn is found to be roughly twice of that for Xe across the range of studied wavelenghts, while the one of Og is much larger still, with its magnitude being 4.5 times larger for the longest to about 6.5 times larger for the shortest wavelenghts. This is in stark contrast to what is observed for the spin-free calculation presented in Table 5, where we see changes of only a factor of roughly 1.5 going from Xe to Rn, and of 1.8 going from Rn to Og. This is consistent with prior theoretical investigations of Og, that predict this element to be much more polarizable due to the extremely large spin-orbit effects in this superheavy element Jerabek2019a, Jerabek2019b.
While the first excited state is lowered due to relativity, this strongly increased polarizability due to spin-orbit effects manifests itself already far away from this resonance. To document this, we present in LABEL:SI-tab:excitation-ng calculations for the lowest excited states of these atoms. These show that while the first excited state is progressively lowered by spin-orbit coupling as the atom gets heavier, the wavelengths probed are still far away from the resonances.
@@
To further complete the picture, we computed the 2DCM dipole polarizabilities for the noble gases with LR-CC and LR-EOMCC for the same range of frequencies as for the Verdet constant. The results, presented in LABEL:SI-tab:polar-ng-table, are in line with the findings for the static polarizabilties for the HX systems and our work Yuan_Xiang_LR in that there is no appreciable difference between the two CC variants for this property.
Our results are in good agreement with the experimental static dipole polarizabilities for Xe and Rn (27.3 and 35.8 a.u. respectively) Hohm1994, CRC2014, and likewise for the theoretical predictions for Og (57.98 a.u.) of Smits and coworkers Smits2023. Though we did not calculated values for atomic Ts, our results from HTs capture well the trend of decreasing polarizability from Ts to Og (76.3 to 57.98 a.u.) Smits2023 predicted in that work. We attribute the observed discrepancies to the lack of higher excitations in CCSD.
@@
To check for the contribution of subvalence electrons, we also considered additional calculations in which we retained the virtual space truncation at 5 a.u. but included also the electrons. In our prior QR-EOMCC calculations Yuan_Xiang_QR we found, based on non-relativistic calculations (the difference of calculations with the same correlating space, and by correlating all occupied and virtual orbitals), that the effect of correlating electrons beyond the valence shell for the Xe atom could result in a decrease of the Verdet constant by about -0.08, -0.19 and -0.28 [10-3 rad/(T m)], for 1064 nm, 694.3 nm and 589 nm respectively.
Our current results, shown in LABEL:SI-tab:verdet-table-scan-n-1dnsnp, confirm that a decrease in Verdet constant is indeed observed for all atoms, though with the current virtual space truncation for Xe we recover slightly smaller changes (-0.07, -0,18 and -0.26 [10-3 rad/(T m)], for 1064 nm, 694.3 nm and 589 nm respectively). This is consistent with our earlier finding, that upon correlating the electrons, the excitation energies for all systems are (slightly) shifted upward. Another change seen upon correlating the electrons is that the difference between QR-CC and QR-EOMCC values increases somewhat, though in relative terms these differences remain close to 1% overall.
As a first comparison beyond atomic systems, in Table 6 we present the Verdet constants obtained with the 2DCM Hamiltonian for the hydrogen halides up to HI. The values of the unique response functions used to calculate the Verdet constant are given in the SI. whereas systems.
Due to their axial symmetry, for the HX systems we require at least two response function evaluations; in addition to we calculated (and verified that we obtain that it yields the same result as ), so that the Verdet constants are calculated with the expression
| V() | |||||
| System | CC | EOM | % | ||
| HF | 0.01 | 0.0161 | 0.0162 | -0.0001 | -0.41 |
| 0.02 | 0.0647 | 0.0650 | -0.0003 | -0.41 | |
| 0.03 | 0.1461 | 0.1467 | -0.0006 | -0.41 | |
| 0.05 | 0.4112 | 0.4128 | -0.0017 | -0.40 | |
| 0.09 | 1.3954 | 1.4009 | -0.0056 | -0.40 | |
| HCl | 0.01 | 0.0781 | 0.0789 | -0.0008 | -0.99 |
| 0.02 | 0.3133 | 0.3165 | -0.0031 | -0.99 | |
| 0.03 | 0.7089 | 0.7159 | -0.0070 | -0.99 | |
| 0.05 | 2.0042 | 2.0239 | -0.0197 | -0.98 | |
| 0.09 | 6.9238 | 6.9903 | -0.0665 | -0.96 | |
| HBr | 0.01 | 0.1356 | 0.1370 | -0.0014 | -1.07 |
| 0.02 | 0.5445 | 0.5503 | -0.0058 | -1.06 | |
| 0.03 | 1.2335 | 1.2466 | -0.0131 | -1.06 | |
| 0.05 | 3.5036 | 3.5405 | -0.0369 | -1.05 | |
| 0.09 | 12.3195 | 12.4447 | -0.1252 | -1.02 | |
| HI | 0.01 | 0.2689 | 0.2726 | -0.0037 | -1.37 |
| 0.02 | 1.0813 | 1.0960 | -0.0147 | -1.36 | |
| 0.03 | 2.4545 | 2.4879 | -0.0334 | -1.36 | |
| 0.05 | 7.0169 | 7.1115 | -0.0946 | -1.35 | |
| 0.09 | 25.3084 | 25.6359 | -0.3275 | -1.29 | |
Following the trends observed for the atoms, the Verdet constants increase across the HF to HI series. This is in line with the increase in molecular polarizabilities for the series, and a reflex of the Verdet constant representing the change in dynamic polarizability in the presence of the external static magnetic field. Likewise, we have that QR-CC and QR-EOMCC yield very similar results for a given frequency and molecular system. In relative terms, we see that the differences are on par with those for the noble gases, with the exception of HF which shows even smaller differences. We note that as in the case of SHG and the Pockels effect, we have a small decrease in the difference between QR-CC and QR-EOMCC for larger perturbing frequencies.
In view of the arguments put forward in the analysis of the hyperpolarizabilities, and gven the similarity between CC and EOMCC polarizabilities for these systems–and consequently of their ground to excited transition matrix elements–the good agreement for the the Verdet constants suggests the terms connected to the state to state transition moments ( in subsection 2.1) should also be similar.
In order to assess this hypothesis, we have analyzed the contributions from different contractions to the response functions in the Verdet constant case as well. The breakdown of the contributions are found in LABEL:SI-tab:breakdown-qrf-cc-verdet for QR-CC and in LABEL:SI-tab:breakdown-qrf-eomcc-verdet for QR-EOMCC. From them, we observe that the contributions arise essentially from the term across the series, with an increased contribution from the for QR-CC seen for HI. We observe that contributions to these terms are essentially made up of singles for all species, a behavior which is distinct from the hyperpolarizabilities.
A comparison between the trends across the halide series and those for the Pockels effect is also interesting, as it highlight the difference between the perturbing operators : for Pockels, the driver for the difference between QR-CC and QR-EOMCC is the operator, which drives the response of the polarizability along the molecular axis. For verdet, on the other hand, the magnetic operator induces an action perpendicular to the molecular axis.
To conclude this section, we consider that through the analysis of the quadratic response functions, we have been able to identify the mechanisms through which the differences between QR-CC and QR-EOMCC arise for the systems under consideration. We also show that depending on the property under considreration, QR-EOMCC can offer a very similar quality than QR-CC–and as shown previously, while outperforming density functional approximations such as B3LYP Yuan_Xiang_QR–at a substantially less cost.
@@ 5 Conclusions
In this work we have presented an implementation in the GPU-accelerated ExaCorr module of coupled cluster quadratic response (QR-CC) theory which can be used in combination with both (2- and 4-component based) relativistic and non-relativistic Hamiltonians, thus extending our prior implementation quadratic response based on the equation of motion (QR-EOMCC) formalism Yuan_Xiang_QR.
@@
We have showcased this implementation by carrying out a detailed comparison between QR-CC and QR-EOMCC in the calculation of both static and dynamic first hyperpolarizabilities (where we determined dispersion curves for second harmonic generation and the Pockels electro-optical effect) for the hydrogen halide series (for HF to HTs), as well as the calculation of the Verdet constant for noble gases (Xe, Rn and Og) and hydrogen halides (HF to HI). In addition to the comparison of non-relativistic and X2C Hamiltonians as done previously, we have carried out calculations with the X2C molecular mean field approach, and gauged the contribution of spin-orbit coupled and scalar relativistic effects.
For the first hyperpolarizabilities, we find that for the HF molecule the difference between QR-CC and EOM-CC is relatively small (around 4%), in line with the studies reported in the literature employing non-relativistic approaches that do not see significant differences between the two approaches. The differences between QR-CC and QR-EOMCC do however increase as one goes down the periodic table, with differences of about 20% already for HCl and a sign change for in HI. Our findings also point to the fact that such differences are significantly smaller when basis sets without diffuse functions are employed, and are particularly small (from HF to HBr) when double zeta bases such as cc-pVDZ are used. We have attributed this difference to differences in the treatment of terms related to state to state transition moments between QR-CC and QR-EOMCC, based on the analysis of the dominant contributions to the hyperpolarizabilties, and the very good agreement between CC and EOMCC dipole polarizabilities. That said, the overall shape of the dispersion curves remains unchanged, indicating that in qualitative terms the QR-EOMCC model still behaves quite similarly to QR-CC.
The contribution of spin-orbit coupling to the hyperpolarizability is rather small up to HBr, but becomes significant from HI onwards, also for the static case, but the fact that we observe similar trends across the HX series for all Hamiltonians considered indicates that, in spite of the non-additivity between relativisic effects and electron correlation, the differences between QR-CC and QR-EOMCC are driven by electron correlation.
Truncation of the correlating space appears to affect QR-EOMCC more than QR-CC for the component for HBr, but the additional electron correlation recovered does not modify the trends upon going down the series from HF to HBr. We note, however, that results can only be considered converged with respect to virtual space truncation if thresholds are set to higher energies (around 100 a.u. in our tests). Furthermore, core(-valence) as a non-negligible effect on the hyperpolarizability of HBr, calling for more systematic studies. We also found orbital relaxation to be small but non-negligible for the polarizability of HBr (1.7%), and it would be of interest to consider such effect for higher-order properties in future work.
@@
In contrast to the hyperpolarizabilies, for the Verdet constant the difference between QR-CC and EOM-CC is minor and roughly constant percentage-wise (about 1%) for all systems studied. Spin-orbit coupling accounts for a relatively small part of the response (about 4%) in Xe, becomes important for Rn (about 35-40%) and absolutely dominates for Og. For the noble gases, correlating the d electrons in addition to the valence reduces the magnitude of the Verdet constants by about 10%. The comparison of the Verdet constants to the Pockels effect for the hydrogen halides further underscored our argument that the differences between QR-CC and QR-EOMCC are likely driven by the description of state to state transition moments for the operators representing the external static fields (magnetic dipole for Verdet, electric dipole for Pockels). We intend to revisit this point in future work, once our implementation allows for the calculation of excited state properties that may shed further light on these differences.
@@
The calculations presented here, albeit restricted to a relatively small set of atomic and molecular systems serve to highlight similarities, but also sometimes important differences, in behavior between the QR-CC and QR-EOMCC approaches across the periodic table and for different molecular properties. Given the rather favorable computational cost and implementation complexity of QR-EOMCC with respect to QR-CC, it is important to further characterize differences between the two approaches in the calculation of nonlinear properties for heavier element systems.
@@
This research used resources of the Oak Ridge Leadership Computing Facility, which is a DOE Office of Science User Facility supported under Contract DE-AC05-00OR22725 (allocations CHM160, CHM191 and CHP109). XY, LH and ASPG acknowledge funding from projects CPER WaveTech, Labex CaPPA (Grant No. ANR-11-LABX-0005-01), ANR CompRIXS (Grant Nos. ANR-19-CE29-0019 and DFG JA 2329/6-1), ANR SCREECHES (Grant Nos. ANR-24-CE29-0904) the I-SITE ULNE project OVERSEE and MESONM International Associated Laboratory (LAI) (Grant No. ANR-16-IDEX-0004), as well support from the French national supercomputing facilities (Grant Nos. DARI A0150801859, A0170801859) and the HPC center at the University of Lille.
@@
The data (input/output) corresponding to the calculations of this paper are available at the Zenodo repositoryzenodo.17215085.
The Supporting Information is available free of charge on the ACS Publications website at DOI: XXX. Working equations for CCSD quadratic response theory; QR-EOMCC equations; Remarks on state to state transition strenghts; Comparison to analytic QR-CC results from other codes; Comparison between analytic QR-CC to finite-field calculations with ExaCorr ; QR-CC and QR-EOMCC and results for different Hamiltonians for HF to HTs systems; Breakdown of the contributions of different terms of the QR-CC and QR-EOMCC response functions for the components, as well as those to obtain the Verdet constant, for HF to HI systems; Basis set convergence study for QR-CC and QR-EOMCC for the components for HF to HBr systems, using the 2DCM Hamiltonian; Dipole moment and polarizabilities for HF to HTs, and Xe to Og systems; Verdet constants for Xe to Og systems; Excitation energies for Xe to Og systems; sample timings for QR-CC and QR-EOMCC response function evaluations.